# Complexity transitions in chaotic quantum systems: Nonstabilizerness, entanglement, and fractal dimension in SYK and random matrix models

### v2

@@ IV Power-Law Random Banded Matrix model
IV.1 Model
The PLRBM model Mirlin et al. (1996) is conceptually similar to the RP model, as it is a random matrix model with suppressed off-diagonal elements. Nevertheless, its physics and phase diagram have distinct features from the RP model. The PLRBM model is a random matrix ensemble whose elements are given by
@@ IV.1 Model
The model is known to behave like a random matrix theory for $\alpha<\frac{1}{2}$ Mirlin et al. (1996); Xu and Rao (2023). In the regime $\frac{1}{2}<\alpha<1$, the eigenvectors depart from the Porter–Thomas distribution and follow a generalized hyperbolic distribution Bogomolny and Sieber (2018b), signaling the presence of a weakly-ergodic regime distinct from that of the RP model. A recent study on the variable-range SYK2 model reveals clear signatures of this non-ergodic transition at $\alpha=\frac{1}{2}$ in the many-body spectral statistics Legramandi et al. (2024). There is an Anderson critical point at $\alpha=1$ Varga and Braun (2000); Mirlin et al. (1996), where the eigenvectors display proper multifractal behavior that has been intensively studied in the literature Evers and Mirlin (2000); Mirlin and Evers (2000); Evers and Mirlin (2008). In the range $1<\alpha<\frac{3}{2}$, the PLRBM enters a localized yet superdiffusive phase, while it becomes fully localized for $\alpha>\frac{3}{2}$.
@@ IV.2 Results
Although the analysis of the PLRBM model is restricted to smaller system sizes and has more sizable fluctuations, it supports what we conjectured for the RP model: the complexity markers assume distinct values across the ergodic-to-non-ergodic transition only in the presence of a genuine fractal (or multifractal) phase. Indeed, within error bars, in the PLRBM model, the transition points identified by all three complexity markers coincide (or almost coincide) with the Anderson critical point, the only transition captured by the fractal dimension in the thermodynamic limit. We do not report here the results for the ground state behavior of the PLRBM model since they qualitatively resemble that of the RP model considered in Sec. III.2: the transitions occur earlier than in the bulk Buijsman et al. (2025), yet no intermediate phase is observed. This is consistent with expectations, as the markers used do not reveal a clear extended non-ergodic regime even in the bulk of the PLRBM model.
V SYK4+Ising model
In this section, we consider a quantum many-body model subject to a chaotic–integrable transition to contrast with the random-matrix behavior. For this, we choose a hybrid model consisting of the SYK model interpolating towards the Ising model. The SYK model, originally introduced as a phenomenological model for non-Fermi liquids Sachdev and Ye (1993), is a quantum mechanical system consisting of $M$ Majorana fermions with random $q$-body interactions. It has accumulated significant attention due to the combination of strong interactions and solvability in the large-$M$ limit Kitaev (2015); Maldacena and Stanford (2016), making it a paradigmatic system to study quantum chaos beyond RMT. Moreover, randomly interacting many-body systems, like the SYK model, are known to behave drastically different from RMT near the edge of their energy spectrum Altland et al. (2024). This makes it particularly interesting to compare the complexity markers for the ground state of this model to those of the RP model of Sec. III.2. Further, the SYK model is a prototypical system in the study of black holes and quantum gravity Sachdev (2015); Cotler et al. (2017) through the AdS/CFT correspondence Maldacena (1998); Witten (1998). Indeed, at low energies, it exhibits an emergent conformal symmetry and is effectively described by a Schwarzian action Maldacena and Stanford (2016), mirroring the behavior of Jackiw–Teitelboim gravity Maldacena et al. (2016b). Since the duality becomes manifest in the infrared, it is of particular interest to understand the ground-state properties of the model in order to assess the stability of the holographic phase under perturbations.
V.1 Model
@@
The Ising Hamiltonian is given by an anti-ferromagnetic ZZ spin chain without magnetic field, $H_{\rm Ising}=\sum_{i=1}^{N-1}g\sigma^{z}_{i}\sigma^{z}_{i+1}$. It can be written in terms of Majorana operators using the Jordan–Wigner transformation, which maps each Majorana operator to a Pauli string consisting of $N$ Pauli matrices García-Álvarez et al. (2017),
@@
In addition, $H_{{\rm SYK}_{4}}$ commutes with the anti-unitary time reversal operator $\mathcal{T}$. The interplay between $(-1)^{\mathrm{F}}$ and $\mathcal{T}$ determines the RMT ensemble and degeneracy of the ground state of the SYK model Li et al. (2017): for $M\bmod 4=2$ the ensemble corresponds to GUE, for $M\bmod 8=0$ the corresponding ensemble is GOE, and for $M\bmod 8=4$ it is GSE  [^5] The Gaussian symplectic ensemble describes the random $L\times L$ Hermitian quaternionic matrices. The numerical results were obtained by representing the GSE matrices as $2L\times 2L$ complex matrices. For consistency, when comparing the SYK4 results falling in the GSE symmetry class with the corresponding Haar-random value, we take an ensemble of dimension $L^{N-1}$ such that the Hilbert space dimension of SYK4 matches with the dimension of the complex GSE representation.. In contrast, $H_{\text{Ising}}$ as given in Eq. (18) anticommutes with $\mathcal{T}$ due to the presence of the imaginary unit in the Hamiltonian, lifting for $\lambda\neq 0$ the symmetry class to GUE for all $M$.
V.2 Results
@@
Moreover, the SRE exhibits a peculiar $M$-dependent behavior at small $\lambda$, which can be attributed to the characteristics of the underlying RMT ensembles Cotler et al. (2017); Kanazawa and Wettig (2017). Specifically, when $M\bmod 8=0$, $\mathcal{M}_{2}$ begins at a lower value, and then rises to a peak at some small nonzero $\lambda$, followed by a monotonic decrease. This behavior can be comprehended in the following way: for finite-size systems, the GOE ensemble yields lower magic than GUE or GSE for Haar-random states (see Eq. (27) and Fig. 9 in Appendix E). As contributions from the Ising Hamiltonian kick in at $\lambda\neq 0$, the symmetry class is lifted from GOE to GUE, leading to a rise of the value for the SRE in the case of $M\bmod 8=0$. In contrast, for $M\bmod 4=2$ and $M\bmod 8=4$, the SRE starts at its maximal value and the addition of the Ising contribution only decreases it.
@@
As we move further away from $\lambda\sim 0$, all three markers start deviating from their ergodic value without exhibiting a long plateau as in the RP model—indicating that the ground state is rather susceptible to perturbations by the Ising contribution (see Fig. 6), unlike the $\lambda\gtrsim 0.5$ regime. Such fragility of the SYK4 ground state resembles recent observations in a related model interpolating between SYK4 and SYK2 Jasser et al. (2025), which implements a crossover from many-body chaos to single-body chaos, while always maintaining volume-law magic in the ground state. In our model, the perturbed Ising ground state is a stabilizer state; therefore, the transition in magic appears sharper, reminiscent of the transition observed in the RP ground state (Fig. 4).
@@
We expect the ensemble-dependent features discussed above to disappear in the $M\to\infty$ limit. Indeed, the $M\bmod 8=0$ peak seems to progressively move to smaller values of $\lambda$ and $\Delta\mathcal{M}_{2}$ progressively decreases if we compare the same ensembles. Nonetheless, such finite-size effects are particularly relevant for characterizing realistic many-body chaotic systems beyond random matrix universality Altland et al. (2024); Legramandi and Talwar (2024). The analysis highlights the utility of the SRE as a diagnostic tool due to its sensitivity to the model’s symmetry structure, distinguishing it from other markers such as the fractal dimension and entanglement entropy.
VI Conclusion
@@
In this work, we have conducted a comparative study of various complexity markers across families of Hamiltonians that interpolate between ergodic, extended non-ergodic, and localized phases. When a fractal region in the parameter space is present, such as for states in the bulk of the Rosenzweig–Porter model spectrum, the different markers identify high and low complexity regimes in overlapping but distinct parameter regions. Both entanglement entropy and stabilizer Rényi entropy undergo a transition within the fractal phase that is not detected by the fractal dimension, indicating the existence of a non-ergodic yet complex regime with respect to these two markers. Moreover, the stabilizer Rényi entropy reveals an additional extended intermediate regime with non-maximal magic, before eventually dropping to low complexity in correspondence to the localization point. The presence of an extended fractal region seems to be crucial to the appearance of new complexity regimes: when such a region is absent, as in the ground state of the Rosenzweig–Porter model and the bulk of the power-law random banded model, the transition points identified by all complexity markers remain consistent. The behavior of the stabilizer Rényi entropy suggests that it can uncover subtle features in the quantum phase diagram that might otherwise go unnoticed using conventional localization markers. It would be interesting to corroborate these findings through analytical studies, e.g., of the Rosenzweig–Porter model akin to those employed for fractal dimensions Kravtsov et al. (2015).
@@
Moreover, our analysis of the ergodic to integrable transition of the SYK4 + Ising model suggests that stabilizer Rényi entropy may be a more sensitive marker to the underlying symmetries of the model, as indicated by its non-monotonic behavior as compared to the fractal dimension and the entanglement entropy. This observation motivates a more systematic investigation of the relationship between non-stabilizerness and symmetry classes in SYK-type models, extending beyond the cases considered in Refs. Li et al. (2017); Kanazawa and Wettig (2017). Additionally, it would be valuable to study other quadratic perturbations of the SYK model to determine whether the complexity markers studied here can consistently capture the observed fragility of ground-state ergodicity. This is especially relevant for assessing the stability of the holographic phase, which has been argued to remain robust under deformations only within a vanishingly small region in the thermodynamic limit Lunkin et al. (2018)
@@
Investigating the presence of non-ergodic but highly complex regimes would not only shed light on the structure of the quantum phase diagram but also help assess the sensitivity of non-stabilizer magic to non-ergodic behavior and complexity transitions. This could be pursued by analyzing many-body systems undergoing a many-body localization transition, where multifractal phases are expected to arise, such as disordered spin systems Torres‐Herrera and Santos (2017); Macé et al. (2019, 2019) and the Bose–Hubbard model Lindinger et al. (2019); Geißler and Pupillo (2020). Localization also plays a significant role in quantum optimization Altshuler et al. (2010); Knysh and Smelyanskiy (2010); Pino and García-Ripoll (2018); Cao et al. (2024); Stark et al. (2011); Cao et al. (2024), whose target problems are often described by spin-glass models Panchenko (2013); Venturelli et al. (2015); Hauke et al. (2020). In these contexts, our work offers a promising perspective for probing complex characteristics of physical systems approaching the many-body localization transition.
@@
The presence of distinct crossover values at finite system sizes is crucial for both classical and quantum simulations, as the feasible techniques are often constrained by available quantum resources. For instance, in noisy intermediate-scale quantum (NISQ) devices, the bottleneck is the ability to build up entanglement across the device Preskill (2012), while in tensor networks, the reachable bond dimension constrains the accessible entanglement entropy Eisert (2013); Orús (2014, 2019); Montangero et al. (2018). In contrast, the efficiency of classical calculations using the Gottesmann–Knill framework Gottesman (1997, 1998a); Aaronson and Gottesman (2004); Nest (2008); Gidney (2021) as well as the cost of broad classes of fault-tolerant quantum computers AI (2023); Katabarwa et al. (2024) is determined by the number of $T$-gates Bravyi et al. (2019); O’Gorman and Campbell (2017), which is related to the magic in a quantum state Howard and Campbell (2017). In this context, our analysis could guide the identification of the most efficient simulation strategies for each regime.
VII Acknowledgments
@@
We thank Julius Mildenberger for valuable discussions regarding the calculation of magic and Xhek Turkeshi for discussions on the analytical expression of Haar-random states. This project has received funding by the European Union under Horizon Europe Programme - Grant Agreement 101080086 - NeQST and under NextGenerationEU via the ICSC – Centro Nazionale di Ricerca in HPC, Big Data and Quantum Computing. This work was supported by the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number UeMO19-5.1. This project has received funding from the European Union - Next Generation EU, Mission 4 Component 2 - CUP E53D23002240006 and the Italian Ministry of University and Research (MUR) through the FARE grant for the project DAVNE (Grant R20PEX7Y3A), was supported by the Provincia Autonoma di Trento, and by Q@TN, the joint lab between the University of Trento, FBK-Fondazione Bruno Kessler, INFN-National Institute for Nuclear Physics and CNR-National Research Council. S.B. acknowledges CINECA for use of HPC resources under Italian SuperComputing Resource Allocation– ISCRA Class C Project No. DeepSYK - HP10CAD1L3.
Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Mobility edge in the energy spectrum
The transition to the localized phase can be energy-dependent, a phenomenon referred to as a mobility edge Basko et al. (2006); Chanda et al. (2020); Guo et al. (2021); Geißler and Pupillo (2020); Wei et al. (2020); Yousefjani and Bayat (2023). For instance, low-energy states can become localized at lower values of the parameter regime compared to states near the middle of the spectrum. Random banded models are known to exhibit such a mobility edge Sodin (2010). A similar trend is observed in our numerical analysis of the RP model (see Fig. 1c). As a result, for certain values of the control parameter, low-energy states may already be localized while high-energy states remain ergodic. Given this sensitivity to the energy window, it is crucial to investigate how different complexity markers respond to variations in the chosen spectral region.
@@
In Fig. 7a1–c1, we analyze consecutive $5\%$ segments of the energy spectrum. Since the spectrum is symmetric on average, we focus only on the lower half. As we move toward the center of the spectrum, the transition curves systematically shift to the right, indicating that the transition point occurs at higher values of the control parameter, consistent with the presence of a mobility edge. Additionally, the curves corresponding to higher-energy windows quickly collapse onto one another, suggesting rapid convergence. This behavior supports our choice of using the central $20\%$ of the spectrum for the bulk calculations in section III.2, as this range already displays stable and converged behavior.

### v3

@@ 4 Power-Law Random Banded Matrix model
4.1 Model
The PLRBM model [51] is conceptually similar to the RP model, as it is a random matrix model with suppressed off-diagonal elements. Nevertheless, its physics and phase diagram have distinct features from the RP model. The PLRBM model is a random matrix ensemble whose elements are given by
@@ 4.1 Model
The model is known to behave like a random matrix theory for $\alpha<\frac{1}{2}$ [51, 135]. In the regime $\frac{1}{2}<\alpha<1$, the eigenvectors depart from the Porter–Thomas distribution and follow a generalized hyperbolic distribution [122], signaling the presence of a weakly-ergodic regime distinct from that of the RP model. A recent study on the variable-range SYK2 model reveals clear signatures of this non-ergodic transition at $\alpha=\frac{1}{2}$ in the many-body spectral statistics [136]. There is an Anderson critical point at $\alpha=1$ [137, 51], where the eigenvectors display proper multifractal behavior that has been intensively studied in the literature [55, 57, 11]. In the range $1<\alpha<\frac{3}{2}$, the PLRBM enters a localized yet superdiffusive phase, while it becomes fully localized for $\alpha>\frac{3}{2}$.
@@ 4.2 Results
Although the analysis of the PLRBM model is restricted to smaller system sizes and has more sizable fluctuations, it supports what we conjectured for the RP model: the complexity markers assume distinct values across the ergodic-to-non-ergodic transition only in the presence of a genuine fractal (or multifractal) phase. Indeed, within error bars, in the PLRBM model, the transition points identified by all three complexity markers coincide (or almost coincide) with the Anderson critical point, the only transition captured by the fractal dimension in the thermodynamic limit. We do not report here the results for the ground state behavior of the PLRBM model since they qualitatively resemble that of the RP model considered in Sec. 3.2: the transitions occur earlier than in the bulk [138], yet no intermediate phase is observed. This is consistent with expectations, as the markers used do not reveal a clear extended non-ergodic regime even in the bulk of the PLRBM model.
5 SYK4+Ising model
In this section, we consider a quantum many-body model subject to a chaotic–integrable transition to contrast with the random-matrix behavior. For this, we choose a hybrid model consisting of the SYK model interpolating towards the Ising model. The SYK model, originally introduced as a phenomenological model for non-Fermi liquids [52], is a quantum mechanical system consisting of $M$ Majorana fermions with random $q$-body interactions. It has accumulated significant attention due to the combination of strong interactions and solvability in the large-$M$ limit [53, 139], making it a paradigmatic system to study quantum chaos beyond RMT. Moreover, randomly interacting many-body systems, like the SYK model, are known to behave drastically different from RMT near the edge of their energy spectrum [140]. This makes it particularly interesting to compare the complexity markers for the ground state of this model to those of the RP model of Sec. 3.2. Further, the SYK model is a prototypical system in the study of black holes and quantum gravity [60, 61] through the AdS/CFT correspondence [141, 142]. Indeed, at low energies, it exhibits an emergent conformal symmetry and is effectively described by a Schwarzian action [139], mirroring the behavior of Jackiw–Teitelboim gravity [143]. Since the duality becomes manifest in the infrared, it is of particular interest to understand the ground-state properties of the model in order to assess the stability of the holographic phase under perturbations.
5.1 Model
@@
The Ising Hamiltonian is given by an anti-ferromagnetic ZZ spin chain without magnetic field, $H_{\rm Ising}=\sum_{i=1}^{N-1}g\sigma^{z}_{i}\sigma^{z}_{i+1}$. It can be written in terms of Majorana operators using the Jordan–Wigner transformation, which maps each Majorana operator to a Pauli string consisting of $N$ Pauli matrices [144],
@@
In addition, $H_{{\rm SYK}_{4}}$ commutes with the anti-unitary time reversal operator $\mathcal{T}$. The interplay between $(-1)^{\mathrm{F}}$ and $\mathcal{T}$ determines the RMT ensemble and degeneracy of the ground state of the SYK model [145]: for $M\bmod 4=2$ the ensemble corresponds to GUE, for $M\bmod 8=0$ the corresponding ensemble is GOE, and for $M\bmod 8=4$ it is GSE  [^5] The Gaussian symplectic ensemble describes the random $L\times L$ Hermitian quaternionic matrices. The numerical results were obtained by representing the GSE matrices as $2L\times 2L$ complex matrices. For consistency, when comparing the SYK4 results falling in the GSE symmetry class with the corresponding Haar-random value, we take an ensemble of dimension $L^{N-1}$ such that the Hilbert space dimension of SYK4 matches with the dimension of the complex GSE representation.. In contrast, $H_{\text{Ising}}$ as given in Eq. (18) anticommutes with $\mathcal{T}$ due to the presence of the imaginary unit in the Hamiltonian, lifting for $\lambda\neq 0$ the symmetry class to GUE for all $M$.
5.2 Results
@@
Moreover, the SRE exhibits a peculiar $M$-dependent behavior at small $\lambda$, which can be attributed to the characteristics of the underlying RMT ensembles [61, 146]. Specifically, when $M\bmod 8=0$, $\mathcal{M}_{2}$ begins at a lower value, and then rises to a peak at some small nonzero $\lambda$, followed by a monotonic decrease. This behavior can be comprehended in the following way: for finite-size systems, the GOE ensemble yields lower magic than GUE or GSE for Haar-random states (see Eq. (27) and Fig. 9 in Appendix E). As contributions from the Ising Hamiltonian kick in at $\lambda\neq 0$, the symmetry class is lifted from GOE to GUE, leading to a rise of the value for the SRE in the case of $M\bmod 8=0$. In contrast, for $M\bmod 4=2$ and $M\bmod 8=4$, the SRE starts at its maximal value and the addition of the Ising contribution only decreases it.
@@
As we move further away from $\lambda\sim 0$, all three markers start deviating from their ergodic value without exhibiting a long plateau as in the RP model—indicating that the ground state is rather susceptible to perturbations by the Ising contribution (see Fig. 6), unlike the $\lambda\gtrsim 0.5$ regime. Such fragility of the SYK4 ground state resembles recent observations in a related model interpolating between SYK4 and SYK2 [87], which implements a crossover from many-body chaos to single-body chaos, while always maintaining volume-law magic in the ground state. In our model, the perturbed Ising ground state is a stabilizer state; therefore, the transition in magic appears sharper, reminiscent of the transition observed in the RP ground state (Fig. 4).
@@
We expect the ensemble-dependent features discussed above to disappear in the $M\to\infty$ limit. Indeed, the $M\bmod 8=0$ peak seems to progressively move to smaller values of $\lambda$ and $\Delta\mathcal{M}_{2}$ progressively decreases if we compare the same ensembles. Nonetheless, such finite-size effects are particularly relevant for characterizing realistic many-body chaotic systems beyond random matrix universality [140, 147]. The analysis highlights the utility of the SRE as a diagnostic tool due to its sensitivity to the model’s symmetry structure, distinguishing it from other markers such as the fractal dimension and entanglement entropy.
6 Conclusion
@@
In this work, we have conducted a comparative study of various complexity markers across families of Hamiltonians that interpolate between ergodic, extended non-ergodic, and localized phases. When a fractal region in the parameter space is present, such as for states in the bulk of the Rosenzweig–Porter model spectrum, the different markers identify high and low complexity regimes in overlapping but distinct parameter regions. Both entanglement entropy and stabilizer Rényi entropy undergo a transition within the fractal phase that is not detected by the fractal dimension, indicating the existence of a non-ergodic yet complex regime with respect to these two markers. Moreover, the stabilizer Rényi entropy reveals an additional extended intermediate regime with non-maximal magic, before eventually dropping to low complexity in correspondence to the localization point. The presence of an extended fractal region seems to be crucial to the appearance of new complexity regimes: when such a region is absent, as in the ground state of the Rosenzweig–Porter model and the bulk of the power-law random banded model, the transition points identified by all complexity markers remain consistent. The behavior of the stabilizer Rényi entropy suggests that it can uncover subtle features in the quantum phase diagram that might otherwise go unnoticed using conventional localization markers. It would be interesting to corroborate these findings through analytical studies, e.g., of the Rosenzweig–Porter model akin to those employed for fractal dimensions [56].
@@
Moreover, our analysis of the ergodic to integrable transition of the SYK4 + Ising model suggests that stabilizer Rényi entropy may be a more sensitive marker to the underlying symmetries of the model, as indicated by its non-monotonic behavior as compared to the fractal dimension and the entanglement entropy. This observation motivates a more systematic investigation of the relationship between non-stabilizerness and symmetry classes in SYK-type models, extending beyond the cases considered in Refs. [145, 146]. Additionally, it would be valuable to study other quadratic perturbations of the SYK model to determine whether the complexity markers studied here can consistently capture the observed fragility of ground-state ergodicity. This is especially relevant for assessing the stability of the holographic phase, which has been argued to remain robust under deformations only within a vanishingly small region in the thermodynamic limit [148]
@@
Investigating the presence of non-ergodic but highly complex regimes would not only shed light on the structure of the quantum phase diagram but also help assess the sensitivity of non-stabilizer magic to non-ergodic behavior and complexity transitions. This could be pursued by analyzing many-body systems undergoing a many-body localization transition, where multifractal phases are expected to arise, such as disordered spin systems [149, 150, 150] and the Bose–Hubbard model [151, 132]. Localization also plays a significant role in quantum optimization [152, 153, 154, 155, 156, 155], whose target problems are often described by spin-glass models [157, 158, 159]. In these contexts, our work offers a promising perspective for probing complex characteristics of physical systems approaching the many-body localization transition.
@@
The presence of distinct crossover values at finite system sizes is crucial for both classical and quantum simulations, as the feasible techniques are often constrained by available quantum resources. For instance, in noisy intermediate-scale quantum (NISQ) devices, the bottleneck is the ability to build up entanglement across the device [160], while in tensor networks, the reachable bond dimension constrains the accessible entanglement entropy [34, 35, 28, 36]. In contrast, the efficiency of classical calculations using the Gottesmann–Knill framework [113, 67, 114, 115, 116] as well as the cost of broad classes of fault-tolerant quantum computers [161, 162] is determined by the number of $T$-gates [38, 163], which is related to the magic in a quantum state [43]. In this context, our analysis could guide the identification of the most efficient simulation strategies for each regime.
Acknowledgments
We thank Julius Mildenberger for valuable discussions regarding the calculation of magic and Xhek Turkeshi for discussions on the analytical expression of Haar-random states.
Funding information
@@ Acknowledgments
This project has received funding by the European Union under Horizon Europe Programme - Grant Agreement 101080086 - NeQST and under NextGenerationEU via the ICSC – Centro Nazionale di Ricerca in HPC, Big Data and Quantum Computing. This work was supported by the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number UeMO19-5.1. This project has received funding from the European Union - Next Generation EU, Mission 4 Component 2 - CUP E53D23002240006 and the Italian Ministry of University and Research (MUR) through the FARE grant for the project DAVNE (Grant R20PEX7Y3A), was supported by the Provincia Autonoma di Trento, and by Q@TN, the joint lab between the University of Trento, FBK-Fondazione Bruno Kessler, INFN-National Institute for Nuclear Physics and CNR-National Research Council. S.B. acknowledges CINECA for use of HPC resources under Italian SuperComputing Resource Allocation– ISCRA Class C Project No. DeepSYK - HP10CAD1L3. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Mobility edge in the energy spectrum
The transition to the localized phase can be energy-dependent, a phenomenon referred to as a mobility edge [4, 130, 131, 132, 133, 134]. For instance, low-energy states can become localized at lower values of the parameter regime compared to states near the middle of the spectrum. Random banded models are known to exhibit such a mobility edge [129]. A similar trend is observed in our numerical analysis of the RP model (see Fig. 1c). As a result, for certain values of the control parameter, low-energy states may already be localized while high-energy states remain ergodic. Given this sensitivity to the energy window, it is crucial to investigate how different complexity markers respond to variations in the chosen spectral region.
@@
In Fig. 7a1–c1, we analyze consecutive $5\%$ segments of the energy spectrum. Since the spectrum is symmetric on average, we focus only on the lower half. As we move toward the center of the spectrum, the transition curves systematically shift to the right, indicating that the transition point occurs at higher values of the control parameter, consistent with the presence of a mobility edge. Additionally, the curves corresponding to higher-energy windows quickly collapse onto one another, suggesting rapid convergence. This behavior supports our choice of using the central $20\%$ of the spectrum for the bulk calculations in section 3.2, as this range already displays stable and converged behavior.